You require a solution containing 85 %v/v ethanol? You have 1350 mL of a 96 % v/v solution of ethanol. How many L of a 70 %v/v ethanol solution do you need to add to it to make the required 85 %v/v ethanol solution (2 decimal places)
Answer #1

Let us take the given solution of ethanol concentration as C1 = 96% v/v and it’s volume as V1 = 1350ml.

The solution to be added is of concentration 70% v/v. Let us take this concentration as C2 and its volume which is unknown yet as V2.

Let us take the concentration of final required solution as Cf which is = 85%v/v.

The volume of final mixture is = volume of first solution + volume of second added solution

which is 1350+V2. Let us take this as Vf.

Now while adding two similar solutions with different concentrations, we have a formula

CfVf = C1V1 + C2V2

This is obtained by the fact that

volume of ethanol in final solution = volume of ethanol in first solution + volume of ethanol in second solution.

which is (Cf*Vf)100 = (C1V1)100 + (C2V2)100

which gives the above formula CfVf = C1V1 + C2V2

Now substituting the values gives us

85 * (1350 + V2) = (96*1350) + (70 * V2)

(85*1350) + 85V2 = (96*1350) + 70V2

85V2 - 70V2 = (96*1350)-(85*1350)

15V2 = 11*1350

V2 = (11*1350)/15

V2 = 11*90 which is 990 ml or 0.99 litres

The volume of 70% v/v ethanol to be added is 0.99 litres.

Hope this helps you out. Please feel free to give a like. Thank you and have a great day.